### v1

@@ 4 Calibrating Data Generation for Safety-aware Fidelity
Let $\mathcal{D}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{(x,\mathcal{G}(sd))\}$ be the calibration set, containing pairs of real and synthetic data points, where for each pair $(x,\mathcal{G}(sd))$, the real data point $x$ has the scenario description being $sd$, i.e., $x\in\emph{{{{RW}}}}(sd)$. Eq. (6) formulates the design of the calibrator function as a minimization problem
$\emph{{{{argmin}}}}_{\theta}\sum_{(x,\mathcal{G}(sd))\in\mathcal{D}}\emph{{{{loss}}}}(\emph{{{{sia}}}}(\emph{{{{inter}}}}(x,f,sd)),\emph{{{{sia}}}}(\emph{{{{inter}}}}(\mathcal{C}_{\theta}(\mathcal{G}(sd)),f,sd)))$ (6)
where ${\textsf{{loss}}}(x,y)$ is a function computing the difference between $x$ and $y$.
In Eq. (6), given a synthetic data point $\mathcal{G}(sd)$, $\mathcal{C}_{\theta}(\mathcal{G}(sd))$ applies the transformation to create a modified data point. One can then compare the real data point $x$ and the modified synthetic data point $\mathcal{C}_{\theta}(\mathcal{G}(sd))$ in terms of output safety similarity (cf Def. 4). Finally, the design of function ${\textsf{{loss}}}(\cdot,\cdot)$ used in Eq. (6) is subject to user intention. One option is to use ${\textsf{{neq}}}(\cdot,\cdot)$ formally defined in Eq. (7): It evaluates instance-level SA-fidelity in a qualitative/Boolean manner, by setting the value to $1$ if output safety similarity is violated, and 00 if output safety similarity holds. Another option is to use quantitative $L_{1}$ loss following the definition of Eq. (8), summing up the absolute difference of semantic attributes. Other options, such as $L_{2}$ or $L_{\infty}$ loss, can also be used.
${\textsf{{neq}}}(x,y)\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\begin{cases}1&\text{if }x\neq y,\\ 0&\text{otherwise.}\end{cases}$ (7)
${\textsf{{diff}}}_{L1}(\{(a_{1},v_{1}),\ldots,(a_{k},v_{k})\},\{(a_{1},v^{\prime}_{1}),\ldots,(a_{k},v^{\prime}_{k})\})\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\sum^{k}_{i=1}|v_{i}-v^{\prime}_{i}|$ (8)
5 Evaluation
5.1 Research questions
@@
Given a set of simulators/GAI-models, one normally wants to select the best solution for synthetic data generation, considering the computational budget in practice. Thus, it is important to examine whether the proposed SA-fidelity metric is capable of evaluating and comparing these generators with safety awareness.
@@
5.2 Experiment setup
Figure 2: Examples of synthetic images generated from a real-world KITTI image (a), via “style transfer” (b-d) and “scenario understanding” (e).
5.2.1 Synthetic data generation
@@
  • (“Scenario understanding”) Towards accessing RQ3, we additionally build a pipeline to query ChatGPT [^3] In our evaluation, we used ChatGPT version 4: https://chat.openai.com/ to generate scenario descriptions (first as a paragraph, followed by summarizing into the form like Fig. 2(f)) on images from KITTI; then use DALLE-3  [^4] DALLE-3: https://openai.com/dall-e-3 to synthesize new images according to the scenario description. Then manually label the classes and bounding boxes for objects in generated images. This results in the Syn_DALLE3 dataset.
5.2.2 DNNs under testing
@@ 5.3.1 RQ1
In Fig. 4, we present the statistical results of inconsistent predictions, across 3 DNNs (one per column) and 3 synthetic datasets (indicated by boxes of 3 colors). In general, Syn_Carla is observed to have the highest SA-fidelity, while Syn_CN_Depth appears to be the worst. Given such data, statistical hypothesis testing (e.g., T-test or U-test, depending on if the data distribution meets certain assumptions) can be employed to determine if one synthetic dataset is significantly better than another in a principled way. We omit it for simplicity.
Figure 4: Statistics of inconsistent predictions, over 3 DNNs & 3 synthetic datasets.
5.3.2 RQ2
The first step is to identify configurable parameters $\theta$ in $\mathcal{C}_{\theta}$. For image-based inputs, apart from implementing $\mathcal{C}_{\theta}$ also as a DNN, one can also consider simpler implementations such as applying filters from the Python Imaging Library555 https://pillow.readthedocs.io/en/stable/. In this case, $\theta$ can be parameters such as the degree of contrast, sharpness, brightness, or Gaussian blur to be additionally applied once when an image is generated. In RQ2, we adopt such a simple yet sufficient solution to conceptually validate the idea of calibrating $g_{\theta}$. Specifically, we conduct a grid search in the parametric space of 3 parameters representing contrast, sharpness, and brightness of filters. Each parameter ranges from $0.8$ to $1.2$ with a step size of $0.1$.
The results of RQ2 are presented in Table 1. Despite we only searched for the optimal configuration in a very limited parametric space, we may already observe that: (1) The $\mathcal{C}_{\theta}$ indeed can be configured, while the best and worst configurations exhibit significant differences for SA-fidelity; (2) The optimal configuration of a given synthetic data generator varies across different DNNs; (3) The best configuration for SA-fidelity is not necessarily the same as for OV-fidelity, nor is the worst configuration.
SA-Fidelity OV-Fidelity
@@ 5.3.2 RQ2
Table 1: Best and worst configurations for SA-fidelity and OV-fidelity, given a DNN and synthetic data generator. NB, for each entry, the 3 parameters in bracket represent “(contrast, brightness, sharpness)” and the number after colon is the total number of inconsistent predictions over the calibration dataset.
5.3.3 RQ3
@@
6 Related Work
Simulation-based testing for autonomous systems has been an active field with accessible commercial tools. Testing non-perception modules is relatively mature (see [23] for a review), where the scenario description can be translated into objects (as bounding boxes) and their physical characteristics (e.g., velocity, coordinate), serving as the input to the motion planner. Our paper primarily focuses on image-based perception; simulation tools such as NVIDIA DriveSim [14] can create “photo-realistic images” as the underlying pipelines supporting physics-based rendering. These efforts lead to synthetic datasets like SynPeDS [17] tailored for pedestrian detection tasks in urban scenes. With a virtual world modeled, varying scene parameters can reveal performance-limiting factors in perception models [6]. Apart from physic-based synthetic data generation, the use of generative AI also hints at strong promises [24, 8]. Despite the above-mentioned progress being made, fundamental concerns remain regarding the validity of using virtual testing as a strong safety argument (i.e., to replace “real miles” with “virtual miles” [7]). This motivates us to introduce a novel perspective to existing fidelity taxonomies (e.g., [15]) and mathematically precise the meaning of fidelity by aligning it with its intended use, while SA-fidelity provides a rigorous statement to avoid over-perfecting every detail in simulation like the effort in [20].
Within the ML community, the technique of domain adaptation (see [21] for an extensive survey) allows a DNN to perform well on new input domains via techniques such as training the DNN to encourage domain confusion in the feature space. Practically, one can utilize the concept of domain adaption, leading to training with both real and synthetic data included (e.g., [19] on mixing Cityscapes (real) and GTA5 (synthetic)), thereby encouraging a DNN have consistent performance in both real and synthetic images. Our result of SA-fidelity complements these results, as (1) we enable a safety-oriented, task-specific assessment of the synthetic data (i.e., not on the DNN), and (2) our calibration on the synthetic data generator allows the DNN to remain unmodified, thereby acting on top of any existing domain adaptation scheme.
7 Conclusion

### v2

@@ V Calibrating Data Generation for Safety-aware Fidelity
Let $\mathcal{D}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{(x,\mathcal{G}(sd))\}$ be the calibration set, containing pairs of real and synthetic data points, where for each pair $(x,\mathcal{G}(sd))$, the real data point $x$ has the scenario description being $sd$, i.e., $x\in\emph{{{\small{{RW}}}}}(sd)$. Eq. (6) formulates the design of the calibrator function as a minimization problem
${{\small\textsf{{argmin}}}}_{\theta}\!\!\!\!\!\!\!\!\!\!\!\!\sum_{(x,\mathcal{G}(sd))\in\mathcal{D}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!{{\small\textsf{{loss}}}}({{\small\textsf{{sia}}}}({{\small\textsf{{inter}}}}(x,f,sd)),{{\small\textsf{{sia}}}}({{\small\textsf{{inter}}}}(\mathcal{C}_{\theta}(\mathcal{G}(sd)),f,sd)))$ (6)
where ${{\small\textsf{{loss}}}}(x,y)$ is a function computing the difference between $x$ and $y$.
In Eq. (6), given a synthetic data point $\mathcal{G}(sd)$, $\mathcal{C}_{\theta}(\mathcal{G}(sd))$ applies the transformation to create a modified data point. One can then compare the real data point $x$ and the modified synthetic data point $\mathcal{C}_{\theta}(\mathcal{G}(sd))$ in terms of output safety similarity (cf Def. 4). Finally, the design of function ${{\small\textsf{{loss}}}}(\cdot,\cdot)$ used in Eq. (6) is subject to user intention. One option is to use ${{\small\textsf{{neq}}}}(\cdot,\cdot)$ formally defined in Eq. (7): It evaluates instance-level SA-fidelity in a qualitative/Boolean manner, by setting the value to $1$ if output safety similarity is violated, and 00 if output safety similarity holds. Another option is to use quantitative $L_{1}$ loss following the definition of Eq. (V), summing up the absolute difference of semantic attributes. Other options, such as $L_{2}$ or $L_{\infty}$ loss, can also be used.
${{\small\textsf{{neq}}}}(x,y)\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\begin{cases}1&\text{if }x\neq y,\\ 0&\text{otherwise.}\end{cases}$ (7)
${{\small\textsf{{diff}}}}_{L1}(\{(a_{1},v_{1}),\ldots,(a_{k},v_{k})\},\{(a_{1},v^{\prime}_{1}),\ldots,(a_{k},v^{\prime}_{k})\})$
:=$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\sum^{k}_{i=1}|v_{i}-v^{\prime}_{i}|$ (8)
VI Evaluation
VI-A Research questions
@@
Given a set of simulators/GAI-models, one normally wants to select the best solution for synthetic data generation, considering the computational budget in practice. Thus, it is important to examine whether the proposed SA-fidelity metric is capable of evaluating and comparing these generators.
@@
VI-B Experiment setup
Figure 2: Example of a real-world KITTI image (a), and its synthetic images generated via “style transfer” (b-d) and “scenario understanding” (e).
VI-B1 Synthetic data generation
@@
  • (“Scenario understanding”) Towards accessing RQ3, we additionally build a pipeline to query ChatGPT [^5] In our evaluation, we used ChatGPT version 4. to generate scenario descriptions (first as a paragraph, followed by summarizing into the form like Fig. 2(f)) on images from KITTI; then use DALLE-3  [^6] DALLE-3: https://openai.com/dall-e-3 to synthesize new images according to the scenario description. Then manually label the classes and bounding boxes for objects in generated images. This results in the Syn_DALLE3 dataset.
VI-B2 DNNs under testing
@@ VI-C1 RQ1
Figure 4: Statistics of inconsistent predictions, over 3 DNNs and 3 synthetic datasets.
In Fig. 4, we present the statistical results of inconsistent predictions, across 3 DNNs (one per column) and 3 synthetic datasets (indicated by boxes of 3 colors). In general, Syn_Carla is observed to have the highest SA-fidelity, while Syn_CN_Depth appears to be the worst. Given such data, statistical hypothesis testing (e.g., T-test or U-test, depending on if the data distribution meets certain assumptions) can be employed to determine if one synthetic dataset is indeed significantly better than another in a principled way. We omit it for simplicity.
VI-C2 RQ2
The first step is to identify configurable parameters $\theta$ in $\mathcal{C}_{\theta}$. For image-based inputs, apart from implementing $\mathcal{C}_{\theta}$ also as a DNN, one can also consider simpler implementations such as applying filters from the Python Imaging Library777 https://pillow.readthedocs.io/en/stable/. In this case, $\theta$ can be parameters such as the degree of contrast, sharpness, brightness, or Gaussian blur to be additionally applied once when an image is generated. In RQ2, we adopt such a simple yet sufficient solution to conceptually validate the idea of calibrating $g_{\theta}$. Specifically, we conduct a grid search in the parametric space of 3 parameters representing contrast, sharpness, and brightness of filters. Each parameter ranges from $0.8$ to $1.2$ with a step size of $0.1$.
The results of RQ2 are presented in Table VI-C2. Despite we only searched for the optimal configuration in a very limited parametric space, we may already observe that: (1) The $\mathcal{C}_{\theta}$ indeed can be configured, while the best and worst configurations exhibit significant differences for SA-fidelity; (2) The optimal configuration of a given synthetic data generator varies across different DNNs; (3) The best configuration for SA-fidelity is not necessarily the same as for OV-fidelity, nor is the worst configuration.
TABLE I: Best and worst configurations for SA-fidelity and OV-fidelity, given a DNN and synthetic data generator. NB, for each entry, the 3 parameters in bracket represent “(contrast, brightness, sharpness)” and the number after colon is the total number of inconsistent predictions over the calibration dataset.
{tblr}
cells = c, cell13 = c=2, cell15 = c=2, cell31 = r=3, cell61 = r=3, cell91 = r=3, vline3 = -, hline3,12 = -0.08em, hline6,9 = -dashed, SA-Fidelity OV-Fidelity
@@ VI-C2 RQ2
VI-C3 RQ3
@@
Remark 4 (Generalisation to unseen scenarios)
All 3 RQs are answered based on a “calibration dataset” containing paired real-world and synthetic data. In this dataset, each simulated scenario is either implicitly represented by the corresponding real-world data (for RQ1 and RQ2) or explicitly extracted from the real-world data through “scenario understanding” (for RQ3). A crucial underlying assumption, for practicality, is that both the assessment and optimal configuration of the synthetic data generator, derived from this calibration dataset, can be generalized to unseen or new scenarios (without collected real-word data) that will be simulated in the virtual testing.
While we plan to conduct rigorous experiments by splitting calibration and testing datasets in future works to validate such generalisablity, this first work mainly focus on the mathematical foundation and theoretical framework.
VII Conclusion